SLIT-JET INFRARED LASER SPECTROSCOPY OF RADICALS FORMED IN A LASER VAPORIZATION CLUSTER SOURCE

\underline{R. S. WALTERS}, and D. J. NESBITT, JILA, Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO. 80309.

The first results from a new experiment combining a laser vaporization cluster source and a high resolution infrared laser spectrometer are reported. Small organometallic clusters containing laser-ablated Mg and Al atoms are studied via high resolution infrared absorption spectroscopy in the C-H region. Direct infrared measurements on these metalmolecular complexes and other radicals generated in the ablation cluster source reveal detailed information on electronic and geometric structures and provide significant new insights into the unique metal-ligand chemistry possible inside laser plasmas